OpenStudy (anonymous):

Given a right triangle, one leg is 6 cm and the hypotenuse is 10 cm. Use the Pythagorean theorem to find the length of the other leg

ganeshie8 (ganeshie8):

hypotenuse^2 = leg^2 + leg^2

ganeshie8 (ganeshie8):

thats the pythagorean theorem, simply plugin the given values and solve the unknown..

ganeshie8 (ganeshie8):

10^2 = 6^2 + leg^2

ganeshie8 (ganeshie8):

you need to solve the 'leg'